Suppose the risk-free rate is 2.45% and an analyst assumes a market risk premium of 5.30%. Firm A just paid a dividend of $1.31 per share. The analyst estimates the β of Firm A to be 1.45 and estimates the dividend growth rate to be 4.89% forever. Firm A has 260.00 million shares outstanding. Firm B just paid a dividend of $1.59 per share. The analyst estimates the β of Firm B to be 0.83 and believes that dividends will grow at 2.14% forever. Firm B has 186.00 million shares outstanding. What is the value of Firm A?


Submit
Answer format: Currency: Round to: 2 decimal places.


unanswered
not_submitted
#6
Suppose the risk-free rate is 2.44% and an analyst assumes a market risk premium of 6.11%. Firm A just paid a dividend of $1.32 per share. The analyst estimates the β of Firm A to be 1.29 and estimates the dividend growth rate to be 4.48% forever. Firm A has 297.00 million shares outstanding. Firm B just paid a dividend of $1.91 per share. The analyst estimates the β of Firm B to be 0.78 and believes that dividends will grow at 2.70% forever. Firm B has 190.00 million shares outstanding. What is the value of Firm B?


Submit
Answer format: Currency: Round to: 2 decimal places.
Answer #1

1). According to the CAPM,

Required Return for firm A = Risk-free Rate + [Beta * Market risk premium]

= 2.45% + [1.45 * 5.30%] = 2.45% + 7.685% = 10.135%

Share Price for firm A = [Current Dividend * (1 + g)] / [r - g]

= [$1.31 * (1 + 0.0489)] / [0.10135 - 0.0489]

= $1.37406 / 0.05245 = $26.20

Value of Firm A = Share Price for firm A * Shares Outstanding

= $26.20 * 260 million = $6,811.35 million

6). According to the CAPM,

Required Return for firm B = Risk-free Rate + [Beta * Market risk premium]

= 2.44% + [0.78 * 6.11%] = 2.44% + 4.7658% = 7.2058%

Share Price for firm B = [Current Dividend * (1 + g)] / [r - g]

= [$1.91 * (1 + 0.0270)] / [0.072058 - 0.0270]

= $1.96157 / 0.045058 = $43.53

Value of Firm B = Share Price for firm B * Shares Outstanding

= $43.53 * 190 million = $8,271.52 million